OpenStudy (anonymous):

If the average of 20, 30, 40, 50-x, 10+2x is 87. What is the value of x? A. 265 B. 275 C. 285 D. 295

OpenStudy (uri):

This is in Iba's book :P Right? @ArbabShah

OpenStudy (amistre64):

add em up, divide by now many there are

OpenStudy (anonymous):

hahaaa yess test on 30th :( @uri

OpenStudy (amistre64):

20 + 30 + 40 + 50-x +10+2x = 5(87)

OpenStudy (uri):

Aww,Best of luck for it,I have to apply there next year,I still go through those sums,I came to this and i was confused too. :3 @ArbabShah

OpenStudy (amistre64):

150+x = 5(87) x = 5(87) - 150

OpenStudy (anonymous):

@amistre64 Thankyouuuu:D @uri haha thanks hope i clear it :/ dont worry you'll get in easily you're good at math :)

OpenStudy (uri):

I ain't lol,I'm just good at commerce math,Wait that's what i want to get in Iba,you're not bad :3

OpenStudy (anonymous):

Yes you are na i know :p i suck at math i know :p

OpenStudy (uri):

LOOOOOOOOOL no,You are nice.

OpenStudy (anonymous):

hahah :) your siblings are in iba too right?

OpenStudy (uri):

Yes 2 siblings,and my cousin is applying too like he will be with you on 30th.

OpenStudy (anonymous):

wow genius family mashallah :p

OpenStudy (uri):

Haha,not really,Everyone is genius except me. :(

OpenStudy (anonymous):

yes yes really no seriously you are good :)

OpenStudy (uri):

Thanks :3

OpenStudy (anonymous):

:)